# Schwarzschild-Levi-Civita black hole

### v1

@@ Abstract
Inspired by the geometry of the Ernst black hole which interpolates the static spherically symmetric Schwarzschild black hole and the static axially symmetric magnetic Melvin spacetime, in this Letter, we aim to present a solution to Einstein’s field equation in vacuum to interpolate the Schwarzschild black hole and the Levi-Civita spacetime. Hence, we shall call the solution the Schwarzschild-Levi-Civita black hole. Effectively the solution is a one-parameter non-asymptotically flat black hole that is static, axially symmetric, and singular in its axis of symmetry.
@@ I Introduction
The black hole (1) is a 2-parameter static axially symmetric solution of Einstein-Maxwell equations with an external magnetic field $B$ in the polar direction. The solution reduces into the Schwarzschild black hole in the absence of the magnetic field i.e., $B=0,$ and coincides with Melvin magnetic universe with $m=0$. For these two extreme limits usually the solution (1) is also referred to as the Schwarzschild black hole immersed into a magnetic field or the magnetized Schwarzschild black hole. The Ernst black hole has been widely studied in the literature. One of the very recent studies was done by Senjaya in Senjaya where the exact massless scalar quasibound state in the Ernst black hole background is presented. In Becar Becar et al. studied the quasinormal modes of a charged scalar field in the same background spacetime. Gibbons et al. in Gibbons reconsidered the original rotating Ernst black hole and showed that it asymptotically doesn’t mimic the Melvin metric unless the electric charge of the black hole and the external magnetic field satisfy $Q=jB\left(1+\frac{1}{4}j^{2}B^{4}\right)$ in which $Q$ and $B$ are the electric charge of the black hole and the external magnetic field, respectively, and $j$ is the angular momentum of the Kerr-Newman black hole which was used as a seed solution in the Ernst’ procedure Wild . Furthermore, there are significant details on Ernst’s procedure and the rotating Ernst solution with the electromagnetic fields in Gibbons which makes it essential in studying Ernst’s black hole.
In this current study, we present a new black hole solution that is similar to the Ernst static black hole (1) but instead of interpolating the Schwarzschild black hole and Melvin magnetic universe, it interpolates the Schwarzschild black hole and the Levi-Civita spacetime. Both the Schwarzschild black hole and the Levi-Civita spacetime Levi are solutions to Einstein’s equations in vacuum. Hence, unlike the Ernst black hole, the Schwarzschild-Levi-Civita black hole - this is what we shall call our solution hereafter - is also a solution of Einstein’s field equation in vacuum. In general relativity, except the Schwarzschild black hole and the Levi-Civita spacetime which are static, there exists another solution to Einstein’s equations in vacuum which is known as $\gamma$-metric or Zipoy-Voorhees spacetime Z ; V . All these solutions are very well-known in the literature due to their exactness and clarity which make them practical. Having an exact solution to Einstein’s field equation in its own right is very important and deserves to be searched for. In the solution we present in this Letter, we observe not only its exactness but also its simplicity and clarity which in our opinion can be of interest and significant in the community of general relativity.
@@ II Field equations and the black hole solution
where $\lambda_{0}$ is an integration constant. Upon considering (13) and (11) in the field equations (4)-(8) they reduce to the following equations
@@
where $m$ is an integration constant related to the mass of the black hole. Finally, in the spherically symmetric coordinates i.e., $\left\{t,r,\theta,\phi\right\},$ the black hole solution is expressed by the line element
@@
$K=\frac{1008m^{2}}{\lambda_{0}^{4}r^{14}\sin^{8}\theta}-\frac{288\left(2+\sin^{2}\theta\right)m}{\lambda_{0}^{4}r^{13}\sin^{10}\theta}+\frac{192}{\lambda_{0}^{4}r^{12}\sin^{12}\theta}.$ (19)
@@
$K=\frac{1008\tilde{m}^{2}\lambda_{0}^{2}}{\left(X+1\right)^{14}\left(1-Y^{2}\right)^{4}}-\frac{288\left(3-Y^{2}\right)\tilde{m}\lambda_{0}^{2}}{\left(X+1\right)^{13}\left(1-Y^{2}\right)^{5}}+\frac{192\lambda_{0}^{2}}{\left(X+1\right)^{12}\left(1-Y^{2}\right)^{6}}.$ (21)
@@
Remarkably, with $\lambda_{0}=1$ (22) is the well-known Levi-Civita spacetime with the general line element given by
$ds^{2}=\rho^{2\sigma\left(\sigma-1\right)}\left(-dt^{2}+d\rho^{2}\right)+\rho^{2\sigma}dz^{2}+\rho^{2\left(1-\sigma\right)}d\phi^{2},$ (23)
where the Levi-Civita constant is $\sigma=2$. Therefore, the black hole solution (18) may be called the Schwarzschild-Levi-Civita black hole as we mentioned before. The asymptotic behavior of the black hole solution (18) is the same as the Levi-Civita metric i.e.,
$\lim_{r\rightarrow\infty}ds^{2}\rightarrow\lambda_{0}^{2}r^{4}\sin^{4}\theta\left(-dt^{2}+dr^{2}+r^{2}d\theta^{2}\right)+\frac{d\phi^{2}}{\lambda_{0}^{2}r^{2}\sin^{2}\theta},$ (24)
which is the same as Eq. (22). Given these behaviors, we interpret the line element (18) to be the Schwarzschild black hole of mass m in the Levi-Civita spacetime.
@@ III The photon orbit
IV Conclusion
In this compact Letter, we introduced a new exact solution to Einstein’s field equation in vacuum. The idea of the existence of such a solution was taken from the Ernst black hole which is an interpolation between the Schwarzschild black hole and the Melvin universe. Our exact solution is effectively a one-parameter black hole that interpolates the Schwarzschild black hole and the Levi-Civita spacetime such that we called it the Schwarzschild-Levi-Civita black hole. It is static and axially symmetric with the singularity at its axis of symmetry. Unlike the Ernst black hole which is the solution of the Einstein-Maxwell field equations the Schwarzschild-Levi-Civita black hole is the solution of the Einstein equations in vacuum. We have studied the existence of the photonsphere and shown that unlike in the Schwarzschild black hole, there exists no photonsphere outside the black hole. This is an indication that this black hole is a non-trivial extension of the Schwarzschild black hole. In this Letter, we haven’t studied the physical properties of the new solution and it is left to a more detailed article in our future work.

### v2

@@ Abstract
Inspired by the geometry of the Ernst black hole which interpolates the static spherically symmetric Schwarzschild black hole and the static axially symmetric magnetic Melvin spacetime, in this paper, we aim to present a solution to Einstein’s field equation in vacuum to interpolate the Schwarzschild black hole and the Levi-Civita spacetime. Hence, we shall call the solution the Schwarzschild-Levi-Civita black hole. Effectively the solution is a two-parameter non-asymptotically flat black hole that is static, axially symmetric, and singular in its axis of symmetry.
@@ I Introduction
The black hole (1) is a 2-parameter static axially symmetric solution of Einstein-Maxwell equations with an external magnetic field $B$ in the polar direction. The solution reduces into the Schwarzschild black hole in the absence of the magnetic field i.e., $B=0,$ and coincides with Melvin magnetic universe with $m=0$. For these two extreme limits usually the solution (1) is also referred to as the Schwarzschild black hole immersed into a magnetic field or the magnetized Schwarzschild black hole. The Ernst black hole has been widely studied in the literature. One of the very recent studies was done by Senjaya in Senjaya where the exact massless scalar quasibound state in the Ernst black hole background is presented. In Becar Becar et al. studied the quasinormal modes of a charged scalar field in the same background spacetime. Gibbons et al. in Gibbons reconsidered the original rotating Ernst black hole and showed that it asymptotically doesn’t mimic the Melvin metric unless the electric charge of the black hole and the external magnetic field satisfy $Q=jB\left(1+\frac{1}{4}j^{2}B^{4}\right)$ in which $Q$ and $B$ are the electric charge of the black hole and the external magnetic field, respectively, and $j$ is the angular momentum of the Kerr-Newman black hole which was used as a seed solution in the Ernst procedure Wild . Furthermore, there are significant details on Ernst’s procedure and the rotating Ernst solution with the electromagnetic fields in Gibbons which makes it essential to study Ernst’s black hole.
In this current study, we present a new black hole solution that is similar to the Ernst static black hole (1) but instead of interpolating the Schwarzschild black hole and Melvin magnetic universe, it interpolates the Schwarzschild black hole and the Levi-Civita spacetime. Both the Schwarzschild black hole and the Levi-Civita spacetime Levi are solutions to Einstein’s equations in vacuum. Hence, unlike the Ernst black hole, the Schwarzschild-Levi-Civita black hole - this is what we shall call our solution hereafter - is also a solution of Einstein’s field equation in vacuum. In general relativity, except for the Schwarzschild black hole and the Levi-Civita spacetime which are static, there exists another solution to Einstein’s equations in vacuum which is known as $\gamma$-metric or Zipoy-Voorhees spacetime Z ; V . All these solutions are very well-known in the literature due to their exactness and clarity which make them practical. Having an exact solution to Einstein’s field equation in its own right is very important and deserves to be searched for. In the solution presented in this paper, we observe not only its exactness but also its simplicity and clarity, which make the study interesting in its own right.
@@ II Field equations and the black hole solution
where $\lambda_{0}$ is an integration constant. As it will be shown in the sequel, $\lambda_{0}$ is related to the linear mass density of the solution. Upon considering (13) and (11) in the field equations (4)-(8) they reduce to the following equations
@@
where $m$ is another integration constant. Finally, in the spherically symmetric coordinates i.e., $\left\{t,r,\theta,\phi\right\},$ the black hole solution is expressed by the line element
@@
$\mathcal{K}=\frac{1008m^{2}}{\lambda_{0}^{4}r^{14}\sin^{8}\theta}-\frac{288\left(2+\sin^{2}\theta\right)m}{\lambda_{0}^{4}r^{13}\sin^{10}\theta}+\frac{192}{\lambda_{0}^{4}r^{12}\sin^{12}\theta}.$ (19)
@@
$\mathcal{K}=\frac{1008\tilde{m}^{2}\lambda_{0}^{2}}{\left(X+1\right)^{14}\left(1-Y^{2}\right)^{4}}-\frac{288\left(3-Y^{2}\right)\tilde{m}\lambda_{0}^{2}}{\left(X+1\right)^{13}\left(1-Y^{2}\right)^{5}}+\frac{192\lambda_{0}^{2}}{\left(X+1\right)^{12}\left(1-Y^{2}\right)^{6}}.$ (21)
@@
Remarkably, (22) is the well-known Levi-Civita spacetime with the general line element given by
$ds^{2}=\rho^{2\sigma\left(\sigma-1\right)}\left(-dt^{2}+d\rho^{2}\right)+\rho^{2\sigma}dz^{2}+\alpha^{2}\rho^{2\left(1-\sigma\right)}d\phi^{2},$ (23)
where the Levi-Civita constant is $\sigma=2$. Therefore, the black hole solution (18) may be called the Schwarzschild-Levi-Civita black hole as we mentioned before. We should add that since the singularity extends to infinity on the axis of symmetry the mass is not finite and therefore an ADM mass is not defined. Hence, $m$ is not the mass of the black hole. Let us express the solution in cylindrical coordinates to obtain a reasonable mass density. Introducing the transformation
$\rho=rH\left(r\right)\sin\theta,$ (24)
and
$z=rH\left(r\right)\cos\theta,$ (25)
where
$\frac{d}{dr}\ln\left(rH\left(r\right)\right)=\frac{1}{r\sqrt{1-\frac{2m}{r}}},$ (26)
the black hole solution transforms into
$ds^{2}=-\frac{\lambda_{0}^{2}\rho^{4}}{2^{4}}\left(1+\frac{m}{\sqrt{\rho^{2}+z^{2}}}\right)^{4}\left(1-\frac{m^{2}}{\rho^{2}+z^{2}}\right)^{2}dt^{2}+\\ \frac{\lambda_{0}^{2}\rho^{4}}{2^{6}}\left(1+\frac{m}{\sqrt{\rho^{2}+z^{2}}}\right)^{12}\left(d\rho^{2}+dz^{2}\right)+\frac{4}{\lambda_{0}^{2}\rho^{2}\left(1+\frac{m}{\sqrt{\rho^{2}+z^{2}}}\right)^{4}}d\phi^{2}.$ (27)
In comparison with the Weyl-Lewis-Papapetrou class of static and axially symmetric solution
$ds^{2}=-e^{2\upsilon}dt^{2}+e^{2\mu}\left(d\rho^{2}+dz^{2}\right)+B^{2}e^{-2\upsilon}\rho^{2}d\phi^{2},$ (28)
in which $\mu,\nu$ and $B$ are functions of $\rho$ and $z.$, one finds
$e^{2\upsilon}=\frac{\lambda_{0}^{2}\rho^{4}}{2^{4}}\left(1+\frac{m}{\sqrt{\rho^{2}+z^{2}}}\right)^{4}\left(1-\frac{m^{2}}{\rho^{2}+z^{2}}\right)^{2},$ (29)
$e^{2\mu}=\frac{\rho^{4}}{2^{6}}\left(1+\frac{m}{\sqrt{\rho^{2}+z^{2}}}\right)^{12},$ (30)
and
$B=\frac{1}{2}\left(1-\frac{m^{2}}{\rho^{2}+z^{2}}\right).$ (31)
Similar to the Weyl class of solutions, the Killing vector fields of the black hole (27) are $\partial_{t}$ and $\partial_{\phi}.$ Furthermore, (27) in the limit $\rho\rightarrow\infty,$ upon rescaling $\rho\rightarrow 2\rho,$ $z\rightarrow 2z$ and $t\rightarrow\frac{1}{\lambda_{0}}t$ yields
$ds^{2}=-\rho^{4}dt^{2}+\rho^{4}\left(d\rho^{2}+dz^{2}\right)+\frac{1}{\lambda_{0}^{2}}\rho^{-2}d\phi^{2},$ (32)
which is the Levi-Civita metric in Weyl coordinates with $\sigma=2$. As the ADM mass of the Levi-Civita metric is not defined, we introduce the Komar linear mass density which is given by $\mu_{K}=\frac{1}{\lambda_{0}}$. For the detail of calculating the Komar linear mass density of the Levi-Civita metric, we refer to Komar . On the other hand since the asymptotic behavior of the black hole solution (27) is the Levi-Civita metric they naturally admit the same Komar mass density. Therefore the black hole solution (27) is powered by a linear mass on its axis of symmetry namely the $z-$axis with a Komar mass linear density equal to its asymptotic Levi-Civita solution i.e., $\mu_{K}=\frac{1}{\lambda_{0}}.$
Finally, due to the asymptotic behavior of the black hole solution (18) that is the Levi-Civita spacetime, we interpret the line element (18) to be the Schwarzschild black hole formed in the Levi-Civita spacetime ambiance.
@@ III The photon orbit
IV Petrov Type classification
Starting from the complex null tetrad basis one-form in the Newman-Penrose (NP) formalism
$\sqrt{2}\mathbf{l}=\lambda_{0}r^{2}\sin^{2}\theta\sqrt{1-\frac{2m}{r}}dt-\frac{1}{\lambda_{0}r\sin\theta}d\phi,$ (39)
$\sqrt{2}\mathbf{n}=\lambda_{0}r^{2}\sin^{2}\theta\sqrt{1-\frac{2m}{r}}dt+\frac{1}{\lambda_{0}r\sin\theta}d\phi,$ (40)
$\sqrt{2}\mathbf{m}=\frac{\lambda_{0}r^{2}\sin^{2}\theta}{\sqrt{1-\frac{2m}{r}}}dr+i\lambda_{0}r^{3}\sin^{2}\theta d\theta,$ (41)
$\sqrt{2}\mathbf{m}^{\ast}=\frac{\lambda_{0}r^{2}\sin^{2}\theta}{\sqrt{1-\frac{2m}{r}}}dr-i\lambda_{0}r^{3}\sin^{2}\theta d\theta$ (42)
and the nonzero Weyl tensor components
$C_{trtr}=-6\lambda_{0}^{2}r\sin^{2}\theta\left(\cos^{2}\theta\left(m-r\right)-m+\frac{r}{3}\right),$ (43)
$C_{trt\theta}=-6\lambda_{0}^{2}r^{3}\sin^{3}\theta\cos\theta\left(1-\frac{2m}{r}\right),$ (44)
$C_{t\theta t\theta}=\lambda_{0}^{2}r^{3}\sin^{2}\theta\left(1-\frac{2m}{r}\right)\left(\cos^{2}\theta\left(9m-6r\right)-9m+4r\right),$ (45)
$C_{t\phi t\phi}=-\left(1-\frac{2m}{r}\right)\frac{\left(2r-3m\sin^{2}\theta\right)}{\lambda_{0}^{2}r^{5}\sin^{4}\theta},$ (46)
$C_{r\theta r\theta}=\frac{\lambda_{0}^{2}r^{3}\sin^{4}\theta\left(2r-3m\sin^{2}\theta\right)}{\left(1-\frac{2m}{r}\right)},$ (47)
$C_{r\phi r\phi}=-\frac{\left(\cos^{2}\theta\left(9m-6r\right)-9m+4r\right)}{\lambda_{0}^{2}r^{5}\sin^{4}\theta\left(1-\frac{2m}{r}\right)},$ (48)
$C_{r\phi\theta\phi}=-\frac{6\cos\theta}{\lambda_{0}^{2}r^{3}\sin^{3}\theta},$ (49)
$C_{\theta\phi\theta\phi}=\frac{2\left(3\cos^{2}\theta\left(m-r\right)-3m+r\right)}{\lambda_{0}^{2}r^{3}\sin^{4}\theta},$ (50)
we calculate the Weyl-NP scalars $\Psi_{i}$ which are expressed by
$\Psi_{0}=C_{\alpha\beta\gamma\delta}l^{\alpha}m^{\beta}l^{\gamma}m^{\delta}=\frac{-3\left(1-\frac{2m}{r}\right)}{\lambda_{0}^{2}r^{3}\sin^{4}\theta},$ (51)
$\Psi_{1}=C_{\alpha\beta\gamma\delta}l^{\alpha}n^{\beta}l^{\gamma}m^{\delta}=\frac{3\cos\theta\sqrt{1-\frac{2m}{r}}}{\lambda_{0}^{2}r^{6}\sin^{5}\theta},$ (52)
$\Psi_{2}=C_{\alpha\beta\gamma\delta}l^{\alpha}m^{\beta}l^{\gamma}m^{\ast\delta}=-\frac{3\left(1-\left(1-\frac{m}{r}\right)\sin^{2}\theta\right)}{\lambda_{0}^{2}r^{6}\sin^{6}\theta},$ (53)
$\Psi_{3}=C_{\alpha\beta\gamma\delta}l^{\alpha}n^{\beta}m^{\ast\gamma}n^{\delta}=\frac{-3\cos\theta\sqrt{1-\frac{2m}{r}}}{\lambda_{0}^{2}r^{12}\sin^{5}\theta},$ (54)
and
$\Psi_{4}=C_{\alpha\beta\gamma\delta}n^{\alpha}m^{\ast\beta}n^{\gamma}m^{\ast\delta}=\frac{-3\left(1-\frac{2m}{r}\right)}{\lambda_{0}^{2}r^{6}\sin^{4}\theta}.$ (55)
Next, the Petrov type of the solution is determined by the number of the distinct roots $z$ of the following equation
$\Psi_{0}+4z\Psi_{1}+6z^{2}\Psi_{2}+4z^{3}\Psi_{3}+z^{4}\Psi_{4}=0.$ (56)
Having all Weyl-NP scalars $\Psi_{i}$ nonzero, our detailed calculation shows that there are 4 distinct roots for the latter equation, indicating that the spacetime is algebraically general, Petrov type I. This is remarkable to know that both the Schwarzschild black hole and the Levi-Civita spacetime are of Petrov type D which can be found in different limits of the Plebanski-Demianski (PD) 7-parameters class of solutions PD1 ; PD2 . This implies that the Schwarzschild-Levi-Civita black hole is a new solution that departed from the PD class of solutions. Finally, the Weyl-NP scalars once more imply the occurrence of the singularity at the axis of symmetry with $\theta=0,\pi$, and $r=0$.
V Conclusion
In this research, we introduced a new exact solution to Einstein’s field equation in vacuum. The idea of the existence of such a solution was taken from the Ernst black hole which is an interpolation between the Schwarzschild black hole and the Melvin universe. Our exact solution is effectively a two-parameter black hole that interpolates the Schwarzschild black hole and the Levi-Civita spacetime such that we called it the Schwarzschild-Levi-Civita black hole. This has to be added that while the new solution is asymptotically the Levi-Civita spacetime, it doesn’t admit the Schwarzschild black hole solution in any limit. It is more like the Schwarzschild black hole is formed in the Levi-Civita spacetime. This is in analogy to the Ernst solution where a Schwarzschild black hole is formed in the Melvin universe. It is static and axially symmetric with the singularity at its axis of symmetry. Unlike the Ernst black hole which is the solution of the Einstein-Maxwell field equations, the Schwarzschild-Levi-Civita black hole is the solution of the Einstein equations in vacuum. We have studied the existence of the photonsphere and shown that unlike in the Schwarzschild black hole, there exists no photonsphere outside the black hole. This is an indication that this black hole is a non-trivial extension of the Schwarzschild black hole. Further investigation on the physical properties of the new solution is open.